OpenStudy (anonymous):

How are the formulas for the circumference of a circle and area of a circle derived?

OpenStudy (anonymous):

@.Sam. ?

OpenStudy (anonymous):

integration

OpenStudy (anonymous):

By Definite integration

OpenStudy (accessdenied):

The circumference is relatively easy; we determined by experiment that the ratio between circumference (distance around) and diameter (distance through) is roughly a constant. So we have: \( \dfrac{C}{d} = k \) \( C = kd \) That constant of proportionality has been given a name, \( \pi \), which we have estimated to great accuracies. Place that in for k, and we have the formula: \( C = \pi d = 2 \pi r \) \(d = 2r \) Then yes, integration would get you from circumference to area. :)

OpenStudy (anonymous):

" http://www.mathopenref.com/circumferencederive.html "